OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -21 and 567, respectively.

OpenStudy (agent0smith):

use\[\Large a_n = a_1 r^{n-1}\] a2 = -21 so: \[\Large -21 = a_1 r^{2-1} \]so\[\large -21 = a_1 r\] a5 = 567 so: \[\Large 567 = a_1 r^{5-1} \]so\[\large 567 = a_1 r^4\]

OpenStudy (agent0smith):

Now solve these two for a1 and r \[\large -21 = a_1 r\]\[\large 567 = a_1 r^4\] use substitution, solve the first equation for a1, then plug that into the second equation.

OpenStudy (anonymous):

7?

OpenStudy (agent0smith):

7 for what? I don't know, haven't worked it out.

OpenStudy (anonymous):

an = 7 • 3n - 1

OpenStudy (agent0smith):

That won't work, it doesn't give a second term of -21

OpenStudy (anonymous):

7 • (-3)n -

OpenStudy (agent0smith):

^that works, r=-3

OpenStudy (anonymous):

Find the sum of the arithmetic sequence. 17, 19, 21, 23, ..., 35

OpenStudy (anonymous):

i gots 79

OpenStudy (agent0smith):

Has to be bigger than 79 from those numbers. Just fill in the missing spots and add them all together 17+19+21+...+35 = ?

OpenStudy (anonymous):

oh 260

OpenStudy (agent0smith):

Yep